OpenStudy (anonymous):

Which of the following is the general form of the equation 5x2 - 20x = 15? x2 - 4x = 3 x2 - 4x - 3 = 0 5x2 - 20x - 15 = 0

OpenStudy (anonymous):

what do u think? :)

OpenStudy (anonymous):

Well I'm reviewing past Algebra units, and when I did this quiz, the answer was B. But I forget how I got to that answer. I just need someone to explain the solution.

OpenStudy (primeralph):

just move the 15 over to the left

OpenStudy (primeralph):

and then divide all by 5

OpenStudy (anonymous):

well\[5x^2 - 20x = 15\]subtract 15 from both sides\[5x^2 - 20x -15= 15-15\]\[5x^2 - 20x -15= 0\]divide by 5\[x^2-4x-3=0\]

OpenStudy (anonymous):

Ohhhhhh. Makes so much sense! Thanks!!!

OpenStudy (primeralph):

nothing for me?

OpenStudy (anonymous):

very welcome :)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

@primeralph I couldn't do both, so I just went with who got to my question first. Sorry. I'll have more questions though! ;)

OpenStudy (primeralph):

@mukushla you really shouldn't be posting answers.

OpenStudy (anonymous):

answers with enough explanations are allowed here dude :)

OpenStudy (primeralph):

No. They said no.

OpenStudy (primeralph):

I don't have a prob;em with it, but that's what they said.

OpenStudy (anonymous):

i think thats for the direct answers...and yes u r right we should guide students through problems and letting them to do part of it, but thats not necessary all the time

OpenStudy (anonymous):

and one more point, when a helper starts to answering someone, we must be patient and let them do their job :)

OpenStudy (primeralph):

If you're referring to this case, you did not start helping. You posed a question.

OpenStudy (anonymous):

@mukushla Agreed.